# Activation-Informed Merging of Large Language Models

### v1

@@ 1 Introduction
Foundation models are rapidly becoming the dominant force for building Artificial Intelligence (AI) systems. In many cases, researchers build their machine learning models by starting from pre-trained foundation models and fine-tuning (FT) these pre-trained models for some desired target task Pareja et al. (2024). In such a paradigm, numerous fine-tuned models are developed for various tasks. However, an important opportunity is missed, as these fine-tuned task-specialized models typically operate in isolation without leveraging the rich features that each possesses Sudalairaj et al. (2024). This fact highlights the importance of a growing area of research focused on combining multiple task-specialized models fine-tuned from the same base foundation model.
In particular, as large language models (LLMs) continue to evolve, it becomes increasingly important to develop methods that can effectively fuse the specialized knowledge of various fine-tuned models derived from the same foundation model. Model merging has shown broad applications, including enhancing accuracy and robustness Wortsman et al. (2022), improving generalization Ramé et al. (2023), multi-modal models Sung et al. (2023), and model alignment to human feedback Rame et al. (2024); Ramé et al. (2024). Given these benefits, a substantial amount of attention has been devoted to developing more effective merging algorithms for LLMs.
In the vast majority of cases, merging LLMs is done using algorithms that explore the weight space of models and do not leverage the information in the activation space. Activation space information has been widely used to develop model pruning and compression methods, both in the context of general deep learning methods Frantar & Alistarh (2022), and more specifically for large language models Lin et al. (2024). However, this direction has remained under-explored for developing more robust merging algorithms. We hypothesize that the activation space information may hold key insights useful for model merging and explore this in our work.
@@
AIM fundamentally relates to the widely used approach of weight regularization in continual learning Kirkpatrick et al. (2017); Ritter et al. (2018); Aljundi et al. (2018). When merging large language models fine-tuned from the same base model, the goal is to maintain the base model’s performance while improving the merged model’s expertise using the fine-tuned models. Various methods have been proposed to merge fine-tuned LLMs Yu et al. (2024b, a); Ilharco et al. (2023); Wortsman et al. (2022); Davari & Belilovsky (2024); Yadav et al. (2023); Matena & Raffel (2022); Jin et al. (2023). Although useful, these methods are fragile to outlier and low-quality fine-tuned models and may perform worse than the base model. Hence, we take a new perspective toward merging and adopt the continual learning view to prevent catastrophic forgetting. Our extensive experimental study shows that AIM is a complementary solution that can be applied in conjunction with any prior merging methods and consistently improves their performances across all tested benchmarks—including math, code, and instruction following— by up to 40%. Despite its simplicity, AIM shows the effectiveness and importance of the activation space information for more effective model merging.
@@
2 Background & Related Work
@@
Fine-tuning pre-trained language models has become the prevalent paradigm for building downstream NLP models. Often, fine-tuned models are readily available, but their training data is not due to data privacy or intellectual property concerns Dodge et al. (2020). With many fine-tuned checkpoints from the same pre-trained model, LLM merging has emerged as a complementary approach to fine-tuning, combining multiple task-specialized models into a unified architecture. This technique offers several advantages: it reduces storage and inference costs by consolidating expertise into a single model, facilitates compositional generalization by integrating specialized capabilities from different fine-tuned models, and supports decentralized model development by allowing independently trained models to be merged efficiently Yadav et al. (2024).
@@
Model soup, proposed by Wortsman et al. (2022), demonstrates even simple averaging of the weights of multiple fine-tuned models can enhance accuracy and robustness in image classification and natural language processing applications. Further, Ramé et al. (2023) show that model merging can improve out-of-distribution generalization performance. Extending this approach beyond unimodal settings, Sung et al. (2023) empirically demonstrate that model merging is also effective in multimodal setups. Beyond generalization benefits, model merging has also been explored for alignment; WARP Ramé et al. (2024) and WARM Rame et al. (2024) introduce weight merging strategies to improve alignment in reinforcement learning from human feedback. WARP demonstrates that merging rewarded policies enhances model quality and alignment, while WARM shows that weight-averaged reward models improve robustness and help mitigate reward hacking.
@@
Recently, many methods have been proposed to go beyond simple weight averaging to merge fine-tuned LLMs into a multitask model by combining their capabilities. Spherical Linear intERPolation (SLERP) Shoemake (1985), originally proposed for animating rotations, which interpolates between two checkpoints, can be seen as a modification to simple weight averaging that finds a spherical path instead of a linear path in models’ parameter space. A main shortcoming of this approach is that it does not support merging more than two models. Jang et al. (2025) also leverages geometric insights, showing that merging only two fine-tuned models can provide superior in-distribution and out-of-distribution performance compared to ensembling multiple models.
@@
Task Arithmetic Ilharco et al. (2023) generalizes simple weight averaging by introducing task vectors. It suggests moving the base model parameters in the direction of a weighted average of fine-tuned model differences with respect to the base model. Followed by the introduction of task vectors, Model Breadcrumbs Davari & Belilovsky (2024), Trim, Elect Sign & Merge (TIES merging)  Yadav et al. (2023), and Drop And REscale (DARE) Yu et al. (2024a) leverage pruning techniques for better and more scalable ways of merging task vectors. WeIght DisENtanglement based merging (WIDEN) Yu et al. (2024b) takes a more sophisticated approach to model merging by disentangling and analyzing weight components.
@@
Another line of work on model merging takes advantage of the information in the model activations of the training data. Matena & Raffel (2022) propose Fisher merging, which leverages the Laplace approximation by using the diagonal of each model’s Fisher information. Jin et al. (2023) attempt to minimize prediction differences between the merged model and the individual models and introduce the Regression Mean (RegMean) method that calculates the optimal weights with respect to Euclidean distance to model predictions.
@@
Continual learning strategies can be categorized into five overarching approaches: regularization, where parameters are constrained or updated based on past data; replay, where past data is replayed for the model as it encounters new data; optimization, where the loss function and optimizer are targeted; representation, where new data representations and learned embeddings can be exploited for less forgetting; and architecture, where models and parameters can expand as new data arrives Wang et al. (2024a). To avoid catastrophic forgetting of the base model’s abilities, we view the model merging problem through the lens of CL, primarily focusing on regularization-based methods. Regularization-based methods penalize deviation from the base model according to some norm Shi et al. (2024). Various methods have been proposed to mitigate catastrophic forgetting such as Aljundi et al. (2018); Kirkpatrick et al. (2017); Ritter et al. (2018). In particular, Elastic Weight Consolidation (EWC) employs a Fisher Information Matrix to identify and protect parameters crucial for previous tasks by adding a quadratic penalty on the parameter shifts. Integrating CL approaches into the merging framework involves defining a weighted regularization term that selectively constrains parameter updates in critical areas for retaining previously learned tasks. This integration not only mitigates the risk of catastrophic forgetting but also enhances the adaptability and utility of the merged model.
@@
Using activation space information has been shown to be useful in the context of model compression. Frantar & Alistarh (2022) show that using a calibration dataset, deep learning models can be quantized and/or pruned efficiently. Lin et al. (2024) introduce Activation-aware Weight Quantization (AWQ) for LLM compression and show that protecting only 1% salient weights can greatly reduce quantization error.
@@ 3.1 The Merging Problem and Connections to Continual Learning
Consider the merging of $N$ models with parameters $\bm{\theta}_{1},\bm{\theta}_{2},\cdots,\bm{\theta}_{N}$ fine-tuned on different tasks from a common pre-trained model with parameters $\bm{\theta}_{pre}$. For each fine-tuned LLM, we are essentially creating experts on specific tasks that move away from the generalist pre-trained model, hence usually degrading performance in some tasks while improving performance on the task for which the model is fine-tuned. In this sense, each FT model with parameters $\bm{\theta}_{n}$ can be seen as a model fitted to a new task $\mathcal{D}_{n}=\{X_{n},Y_{n}\}$ in a continual learning scenario with the potential for catastrophic forgetting on the generalist pre-trained model, which may not perform as well on the specific task but will have a more balanced performance across various tasks. As such, we hypothesize that when merging FT LLMs adapted from the same base model, emphasis on the base model can build better robustness to large performance degradation across numerous tasks while still allowing capturing each FT expert’s capabilities. AIM seeks to achieve this by relaxing the changes to the salient weights of the base model in the final merged model. In this way, AIM is analogous to weight regularization in many continual learning approaches Wang et al. (2024b); Ritter et al. (2018); Kirkpatrick et al. (2017); Aljundi et al. (2018); Shi et al. (2024). Notably, the saliency of weights is determined by analyzing the activation space of the base model rather than just regularizing based on the weight space.
@@ 3.2 Activation Space Analysis
AIM determines the saliency of a model’s weights by looking at the scale of activations by passing a calibration dataset to the model and recording the scale of activations in each channel. To better understand why, we will analyze how the perturbation of weights of a given model affects the model outputs. Let the original weights be $w\in\mathbb{R}^{N\times M}$ and the perturbation be $\delta w\in\mathbb{R}^{N\times M}$, such that the perturbed weights are:
$w^{\prime}=w+\delta w$ (1)
The output of a linear layer with input $x\in\mathbb{R}^{N}$ and perturbed weights is:
@@
The error due to the perturbation is:
$\text{Error}=y^{\prime}-y=x(\delta w)$ (3)
And the magnitude of the error scales with the magnitude of activation $x$:
$\|\text{Error}\|_{p}=\|x\delta w\|_{p}\leq\|x\delta w\|_{1}\leq\sum_{i=1}^{N}\left(|x_{i}|\sum_{j=1}^{M}|\delta w_{ij}|\right)$ (4)
For any specific input channel $x_{i}$, the error contribution from perturbation in the $i$-th row of $w$, $\delta w_{i}$ is amplified by the magnitude of the same channel in the input. As such, one could selectively regularize the weights based on the importance of the input channels, i.e., their magnitudes. In this way, we use a calibration dataset to capture the average magnitude of the input channels for each layer of the base model and determine the saliency of weights in the base model from the activation space. We choose the calibration dataset to be a subset of the validation data from the pile dataset Gao et al. (2020) which is similar in distribution to most pre-training data and is the choice of calibration data for AWQ Lin et al. (2024) as well (based on the latest implementation on GitHub at the time of writing). This calibration dataset is considered to be fairly diverse and not task-specific, while also being similar to most pre-training data, which should allow for capturing a representative set of activation magnitudes, and hence salient weights. Using this data to capture the average magnitudes of the input activation, we come up with a complementary solution to apply relaxation to weights of merged models using any merging method and show that our adaptable relaxation scheme helps improve the performance of merged models significantly.
@@ 3.3 Adaptable Relaxation Scheme
As discussed, we introduce an adaptable relaxation scheme based on the activations of the base model, which we wish not to stray away from significantly. To make the scheme adaptable to any merging algorithm in the weight space, we formulate our relaxation scheme in terms of the changes in the weights. Given a task-agnostic representative calibration corpus $D$. We can pass this corpus through the model and accumulate the activations from each token in the dataset. This will yield the average magnitude of activations for each channel in all layers of the model. Let vector $a_{i}$ be the normalized (i.e., normalized by division by the largest activation magnitude in layer $i$) average of each activation channel in layer $i$ given the calibration corpus $D$. We then associate these activations with the corresponding model parameters, such as the columns/rows of weight matrices in linear layers. This results in a per-weight activation mapping, denoted by $A$, which has the same dimensionality as the model parameters $\bm{\theta}$. Next, we define the action of any given merging method by the changes that it applies to the model weights with respect to each of the fine-tuned models being merged (i.e., $\bm{\theta}_{1},\bm{\theta}_{2},\ldots,\bm{\theta}_{N}$). Specifically, we denote the weight update contributed by a fine-tuned model (with parameters $\bm{\theta}_{i}$) to the model parameters by $\bm{\Delta}_{i}$ (e.g, $\bm{\Delta}_{i}=\frac{\bm{\theta}_{i}-\bm{\theta}_{\text{pre}}}{N}$ for weight averaging). Now, we propose an adaptive relaxations scheme that adjusts the final model as follows:
$\bm{\theta}_{\text{merged}}=\bm{\theta}_{\text{pre}}+(1-|A_{\text{pre}}|(1-\omega))\sum_{i=1}^{N}\lambda_{i}\cdot\bm{\Delta}_{i},$ (5)
@@
where $\omega$ is the relaxation factor that controls how much relaxation is applied (an $\omega$ of 0.0 would revert the most salient weight to the base weights and an $\omega$ of 1.0 applies no relaxation), and $\lambda_{i}$ are the weight factors for each $\Delta_{i}$ and the subscript pre refers to the pre-trained model. In this work, since we do not explicitly look at the merging algorithm’s inner workings, we can simply fuse the terms $\sum_{i=1}^{N}\lambda_{i}\cdot\bm{\Delta}_{i}$ into a single algorithm-agnostic term $\Delta_{\text{merged}}$ and simplify the relaxation scheme to:
$\bm{\theta}_{\text{merged}}=\bm{\theta}_{\text{pre}}+(1-|A_{\text{pre}}|(1-\omega))\Delta_{\text{merged}}.$ (6)
@@
In our experiments, we apply this relaxation scheme to several different merging methods and explore how the hyper-parameter $\omega$ affects the merged model’s behavior, and present these results in Section 5.

### v2

@@ 1 Introduction
Foundation models are rapidly becoming the dominant force for building Artificial Intelligence (AI) systems. In many cases, researchers build their machine learning models by starting from pre-trained foundation models and fine-tuning (FT) these pre-trained models for some desired target task (Pareja et al., 2024). In such a paradigm, numerous fine-tuned models are developed for various tasks. However, an important opportunity is missed, as these fine-tuned task-specialized models typically operate in isolation without leveraging the rich features that each possesses (Sudalairaj et al., 2024). This fact highlights the importance of a growing area of research focused on combining multiple task-specialized models fine-tuned from the same base foundation model.
In particular, as large language models (LLMs) continue to evolve, it becomes increasingly important to develop methods that can effectively fuse the specialized knowledge of various fine-tuned models derived from the same foundation model. Model merging has shown broad applications, including enhancing accuracy and robustness (Wortsman et al., 2022), improving generalization (Ramé et al., 2023), multi-modal models (Sung et al., 2023), and model alignment to human feedback (Rame et al., 2024; Ramé et al., 2024). Given these benefits, a substantial amount of attention has been devoted to developing more effective merging algorithms for LLMs.
In the vast majority of cases, merging LLMs is done using algorithms that explore the weight space of models and do not leverage the information in the activation space. Activation space information has been widely used to develop model pruning and compression methods, both in the context of general deep learning methods (Frantar and Alistarh, 2022), and more specifically for large language models Lin et al. (2024). However, this direction has remained under-explored for developing more robust merging algorithms. We hypothesize that the activation space information may hold key insights useful for model merging and explore this in our work.
@@
AIM fundamentally relates to the widely used approach of weight regularization in continual learning (Kirkpatrick et al., 2017; Ritter et al., 2018; Aljundi et al., 2018). When merging large language models fine-tuned from the same base model, the goal is to maintain the base model’s performance while improving the merged model’s expertise using the fine-tuned models. Various methods have been proposed to merge fine-tuned LLMs (Yu et al., 2024b, a; Ilharco et al., 2023; Wortsman et al., 2022; Davari and Belilovsky, 2024; Yadav et al., 2023; Matena and Raffel, 2022; Jin et al., 2023). Although useful, these methods are fragile to outlier and low-quality fine-tuned models and may perform worse than the base model. Hence, we take a new perspective toward merging and adopt the continual learning view to prevent catastrophic forgetting. Our extensive experimental study shows that AIM is a complementary solution that can be applied in conjunction with any prior merging methods and consistently improves their performances across all tested benchmarks—including math, code, and instruction following— by up to 40%. Despite its simplicity, AIM shows the effectiveness and importance of the activation space information for more effective model merging.
2 Background and Related Work
@@
Fine-tuning pre-trained language models has become the prevalent paradigm for building downstream NLP models. Often, fine-tuned models are readily available, but their training data is not due to data privacy or intellectual property concerns (Dodge et al., 2020). With many fine-tuned checkpoints from the same pre-trained model, LLM merging has emerged as a complementary approach to fine-tuning, combining multiple task-specialized models into a unified architecture. This technique offers several advantages: it reduces storage and inference costs by consolidating expertise into a single model, facilitates compositional generalization by integrating specialized capabilities from different fine-tuned models, and supports decentralized model development by allowing independently trained models to be merged efficiently (Yadav et al., 2024).
@@
Model soup, proposed by Wortsman et al. (2022), demonstrates even simple averaging of the weights of multiple fine-tuned models can enhance accuracy and robustness in image classification and natural language processing applications. Further, Ramé et al. (2023) show that model merging can improve out-of-distribution generalization performance. Extending this approach beyond unimodal settings, Sung et al. (2023) empirically demonstrate that model merging is also effective in multimodal setups. Beyond generalization benefits, model merging has also been explored for alignment; WARP (Ramé et al., 2024) and WARM (Rame et al., 2024) introduce weight merging strategies to improve alignment in reinforcement learning from human feedback. WARP demonstrates that merging rewarded policies enhances model quality and alignment, while WARM shows that weight-averaged reward models improve robustness and help mitigate reward hacking.
@@
Recently, many methods have been proposed to go beyond simple weight averaging to merge fine-tuned LLMs into a multitask model by combining their capabilities. Spherical Linear intERPolation (SLERP) (Shoemake, 1985), originally proposed for animating rotations, which interpolates between two checkpoints, can be seen as a modification to simple weight averaging that finds a spherical path instead of a linear path in models’ parameter space. A main shortcoming of this approach is that it does not support merging more than two models. Jang et al. (2025) also leverages geometric insights, showing that merging only two fine-tuned models can provide superior in-distribution and out-of-distribution performance compared to ensembling multiple models.
@@
Task Arithmetic (Ilharco et al., 2023) generalizes simple weight averaging by introducing task vectors. It suggests moving the base model parameters in the direction of a weighted average of fine-tuned model differences with respect to the base model. Followed by the introduction of task vectors, Model Breadcrumbs (Davari and Belilovsky, 2024), Trim, Elect Sign & Merge (TIES merging) (Yadav et al., 2023), and Drop And REscale (DARE) (Yu et al., 2024a) leverage pruning techniques for better and more scalable ways of merging task vectors. WeIght DisENtanglement based merging (WIDEN) (Yu et al., 2024b) takes a more sophisticated approach to model merging by disentangling and analyzing weight components.
@@
Another line of work on model merging takes advantage of the information in the model activations of the training data. Matena and Raffel (2022) propose Fisher merging, which leverages the Laplace approximation by using the diagonal of each model’s Fisher information. Jin et al. (2023) attempt to minimize prediction differences between the merged model and the individual models and introduce the Regression Mean (RegMean) method that calculates the optimal weights with respect to Euclidean distance to model predictions.
@@
Continual learning strategies can be categorized into five overarching approaches: regularization, where parameters are constrained or updated based on past data; replay, where past data is replayed for the model as it encounters new data; optimization, where the loss function and optimizer are targeted; representation, where new data representations and learned embeddings can be exploited for less forgetting; and architecture, where models and parameters can expand as new data arrives Wang et al. (2024a). To avoid catastrophic forgetting of the base model’s abilities, we view the model merging problem through the lens of CL, primarily focusing on regularization-based methods. Regularization-based methods penalize deviation from the base model according to some norm (Shi et al., 2024). Various methods have been proposed to mitigate catastrophic forgetting such as Aljundi et al. (2018); Kirkpatrick et al. (2017); Ritter et al. (2018). In particular, Elastic Weight Consolidation (EWC) employs a Fisher Information Matrix to identify and protect parameters crucial for previous tasks by adding a quadratic penalty on the parameter shifts. Integrating CL approaches into the merging framework involves defining a weighted regularization term that selectively constrains parameter updates in critical areas for retaining previously learned tasks. This integration not only mitigates the risk of catastrophic forgetting but also enhances the adaptability and utility of the merged model.
@@
Using activation space information has been shown to be useful in the context of model compression. Frantar and Alistarh (2022) show that using a calibration dataset, deep learning models can be quantized and/or pruned efficiently. Lin et al. (2024) introduce Activation-aware Weight Quantization (AWQ) for LLM compression and show that protecting only 1% salient weights can greatly reduce quantization error.
@@ 3.1 The Merging Problem and Connections to Continual Learning
Consider the merging of $N$ models with parameters $\bm{\theta}_{1},\bm{\theta}_{2},\cdots,\bm{\theta}_{N}$ fine-tuned on different tasks from a common pre-trained model with parameters $\bm{\theta}_{\text{pre}}$. For each fine-tuned LLM, we are essentially creating experts on specific tasks that move away from the generalist pre-trained model, hence usually degrading performance in some tasks while improving performance on the task for which the model is fine-tuned. In this sense, each FT model with parameters $\bm{\theta}_{n}$ can be seen as a model fitted to a new task $\mathcal{D}_{n}=\{X_{n},Y_{n}\}$ in a continual learning scenario with the potential for catastrophic forgetting on the generalist pre-trained model, which may not perform as well on the specific task but will have a more balanced performance across various tasks. As such, we hypothesize that when merging FT LLMs adapted from the same base model, emphasis on the base model can build better robustness to large performance degradation across numerous tasks while still allowing for capturing each FT expert’s capabilities. AIM seeks to achieve this by relaxing the changes to the salient weights of the base model in the final merged model. In this way, AIM is analogous to weight regularization in many continual learning approaches Wang et al. (2024b); Ritter et al. (2018); Kirkpatrick et al. (2017); Aljundi et al. (2018); Shi et al. (2024). Notably, the saliency of weights is determined by analyzing the activation space (sensitivities) of the base model rather than just regularizing based on the weight space, similar to Farajtabar et al. (2020); Min et al. (2022)
@@ 3.2 Activation Space Analysis
AIM determines the saliency of a model’s weights by looking at the scale of activations by passing a calibration dataset to the model and recording the scale of activations in each channel. To better understand why, we will analyze how perturbation of the weights of a given model affects the model outputs. Let the original weights be $w\in\mathbb{R}^{N\times M}$ and the perturbation be $\delta w\in\mathbb{R}^{N\times M}$, such that the perturbed weights are $w^{\prime}=w+\delta w$. The output of a linear layer with input $x\in\mathbb{R}^{N}$ and perturbed weights is:
@@
the magnitude of the error due to this perturbation, $\text{Error}=y^{\prime}-y=x(\delta w)$, scales with the magnitude of activation $x$:
$\|\text{Error}\|_{p}=\|\operatorname{diag}(x)\times\delta w\|_{p}\leq\|\operatorname{diag}(x)\times\delta w\|_{1}\leq\sum_{i=1}^{N}|x_{i}|\sum_{j=1}^{M}|\delta w_{ij}|$ (2)
For any specific input channel $x_{i}$, the error contribution from perturbation in the $i$-th row of $w$, $\delta w_{i}$ is amplified by the magnitude of the same channel in the input. As such, one could selectively regularize the weights based on the importance of the input channels, i.e., their magnitudes. In this way, we use a calibration dataset to capture the average magnitude of the input channels for each layer of the base model and determine the saliency of weights in the base model from the activation space.
3.3 Calibration Data & Robustness to Calibration Data
We choose the calibration dataset to be a subset of the validation data from the pile dataset Gao et al. (2020), which is similar in distribution to most pre-training datasets. Most notably, this dataset has been utilized in model compression for both quantization in AWQ Lin et al. (2024) and pruning in WANDA Sun et al. (2024). This calibration dataset is considered to be fairly diverse and not task-specific, which should allow us to quantify weights’ saliency in a robust manner.
Size Doesn’t Matter
The use of activation spaces from calibration data has been thoroughly studied in model compression, by Lin et al. (2024) and Sun et al. (2024). In both cases, authors conduct extensive studies on the robustness of their methods, which, like ours, rely on the magnitude of activations, and both concur in two important findings. 1) When using calibration data to quantify activation space magnitudes to be used by algorithms for model quantization or pruning, performance is robust to dataset quality. This is in contrast to methods that require fine-tuning or retraining after compression. This observation is made clear in both studies and confirms that sensitivity to the quality of the data is much less in methods that require only activation space analysis without training. 2) Most notably, ablation studies on the size of the calibration set also show robustness to the size of calibration data, with both studies confirming that only 8-32 sample sequences of length approx. 2048 tokens are enough for model compression algorithms to produce robust outcomes Lin et al. (2024); Sun et al. (2024). Despite this robustness, we use the same 256 total sequences (approx. 524K tokens) that the authors of both studies use in their main experiments, however, this robustness to dataset size is noted as an important tool for reduction of computational cost and time for running algorithms such as ours. We also study this matter in an ablation study and show that like those studies AIM is also robust to dataset size and can be made much more efficient if need be.
@@ 3.4 Adaptable Relaxation Scheme
As discussed, we introduce an adaptable relaxation scheme based on the activations of the base model, which we wish not to stray away from significantly. To make the scheme adaptable to any merging algorithm in the weight space, we formulate our relaxation scheme in terms of the changes in the weights. Given a task-agnostic representative calibration corpus $D$. We can pass this corpus through the model and accumulate the activations from each token in the dataset. This will yield the average magnitude of activations for each channel in all layers of the model. As previously noted, averaging the inputs over the calibration set yields a vector $x\in\mathbb{R}^{N}$ for each linear layer. Constructing a diagonal matrix from the absolute values of this vector gives $\operatorname{diag}(|x|)\in\mathbb{R}^{N\times N}$. By normalizing $\operatorname{diag}(|x|)$ using its maximum value, we obtain a diagonal matrix $A_{\text{pre}}\in[0,1]^{N\times N}$. This matrix serves to modulate weight updates based on input saliency, allowing for consistent control of merging through a universal hyperparameter $\omega$. Next, we define the action of any given merging method by the changes that it applies to the model weights with respect to each of the fine-tuned models being merged (i.e., $\bm{\theta}_{1},\bm{\theta}_{2},\ldots,\bm{\theta}_{n}$). Specifically, we denote the weight update contributed by a fine-tuned model (with parameters $\bm{\theta}_{i}$) to the model parameters by $\bm{\Delta}_{i}$ (e.g, $\bm{\Delta}_{i}=\frac{\bm{\theta}_{i}-\bm{\theta}_{\text{pre}}}{n}$ for weight averaging). Now, we propose an adaptive relaxation scheme that adjusts the final model. For simplicity in notation let $\bm{\theta}$ refer to weights of a linear layer (i.e. an $N\times M$ matrix), then the AIM relaxation scheme can be written as:
$\delta w_{\text{AIM}}=\bm{\theta}_{\text{merged}}-\bm{\theta}_{\text{pre}}=(1-A_{\text{pre}}(1-\omega))\sum_{i=1}^{N}\lambda_{i}\cdot\bm{\Delta}_{i},$ (3)
@@
where $\delta w_{AIM}$ is the relaxation change and the subscript pre refers to the pre-trained model and $\omega$ is the relaxation factor that controls how much relaxation is applied (an $\omega$ of 0.0 would revert the most salient weight to the base weights and an $\omega$ of 1.0 applies no relaxation), effectively $\omega$ is scaling error/deviation from the base model, and $\lambda_{i}$ are the weight factors for each $\Delta_{i}$. Note that $\lambda_{i}$ is internal to each merging algorithm and not part of the AIM relaxation; however, in general, one could selectively apply this if desired. In this work, since we do not explicitly look at the merging algorithm’s inner workings, we can simply fuse the terms $\sum_{i=1}^{N}\lambda_{i}\cdot\bm{\Delta}_{i}$ into a single algorithm-agnostic term $\Delta_{\text{merged}}$ and simplify the relaxation scheme to:
$\bm{\theta}_{\text{AIM}}=\bm{\theta}_{\text{pre}}+(1-A_{\text{pre}}(1-\omega))\Delta_{\text{merged}}.$ (4)
@@
Note that in general $A_{\text{pre}}$ is not a single matrix rather a mapping of activations for all model parameters obtained as we described prior. In our experiments, we apply this relaxation scheme to several different merging methods and explore how the hyperparameter $\omega$ affects the merged model’s behavior, and present these results in Section 5.
Sensitivity-Based Formulation
An alternative way of choosing the importance scores for changing the model weights—which has been used, e.g., in continual learning Farajtabar et al. (2020); Min et al. (2022), out-of-distribution detection Sharma et al. (2021), and meta-learning Almecija et al. (2022)—is to use the partial derivatives of the (base) model with respect to the parameters, i.e., sensitivities, which correspond to a scaled version of the activations. This is in contrast with using the activations directly, common in model compression. Formally, let $f_{\bm{\theta}}$ denote the model with parameters $\bm{\theta}$, and $\theta$ be one of the weights; then $g_{\theta}=|\frac{\partial f}{\partial\theta}|$ determines how sensitive the output is to perturbing the weight $\theta$. We develop a sensitivity-based formulation of AIM by incorporating sensitivity scores calculated from the gradients and replacing activations with sensitivity scores. We consider $f_{\bm{\theta}}(x)$ to be the logits of the model with parameter $\theta$ and $\mathcal{L}$ to be the entropy function, similar to Farajtabar et al. (2020)’s suggested approach for classification problems. Now let $G_{\text{pre}}=|\frac{\partial\mathcal{L}}{\partial\bm{\theta}}|$ be the magnitude of gradients for all parameters of the pretrained model for samples in a calibration corpus $D$. Then the sensitivity-based formulation can be written as follows:
$\bm{\theta}_{\text{AIM},\text{G}}=\bm{\theta}_{\text{pre}}+(1-G_{\text{pre}}(1-\omega))\Delta_{\text{merged}}.$ (5)
In other words, the sensitivity-based formulation of the problem suggests that regularization of weights through relaxation should be done based on model gradients. However, we note that computing gradients for a calibration corpus can be computationally expensive and will require significantly more memory and compute resources than storing activations, which is akin to performing inference on the model. To verify that using activations retains the same performance and fidelity as this formulation, we conduct our experiments for both AIM and the sensitivity-based formulation. We present the results of the continual learning view in Appendix C, and we see that when comparing results of AIM (Table 1) and the sensitivity-based formulation (Table 3) the performance boost of both approaches is very similar, while AIM being computationally more efficient.